OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3. y - 6 = -3x - 3 y - 6 = -3(x + 1) y = -3x + 3 3x + y = 3

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &(-1\quad ,&6) \end{array} \\\quad \\ slope = m= -3 \\ \quad \\ y-y_1=m(x-x_1)\)

OpenStudy (anonymous):

Thats none of the choices thoo.. @jdoe0001

OpenStudy (anonymous):

your supposed to plot the numbers in

OpenStudy (anonymous):

Oh thanks (: @armine

OpenStudy (anonymous):

i know i did not help lol

OpenStudy (anonymous):

You just did !

OpenStudy (anonymous):

@agent0smith do u know this ???

OpenStudy (agent0smith):

Follow @jdoe0001's instructions... you literally just have to plug in three numbers, and he even gave them to you.